OpenStudy (anonymous):

the length of one side of a square pond is 15 feet. the pond is surrounded by a 3 foot wide walkway. What is the totaly area of the pond and the walkway, in square feet. @Geometry

OpenStudy (anonymous):

area of pond=15*15=225 square feet. area of walkway+pond=21*21=441 square feet

OpenStudy (anonymous):

Thank you! But how did you get 21? And why do you multiply 15 and 15?

OpenStudy (anonymous):

What I mean is why do you multiply 15*15 if you know the answer will be whatever 21*21 is?

OpenStudy (asnaseer):

 I hope the diagram explains it.